The length of a rectangle is 3 inches longer than twice its width w:Write a polynomial expression in standard form in terms of w for the area of the rectangle. Find the area of the rectangle for w=10
Help with this question would be much obliged.

Answer by checkley77(12844)

W=10
L=2*10+3
L*W=AREA
(20+3)*10=AREA
23*10=AREA
230 INCHES IS THE AREA.